# Predictive Coherence and the Moment Hierarchy: Martingale Posteriors for Exchangeable Bernoulli Sequences

### 05 Mar 2026 v2

@@ 11 Optimal stopping and dynamic amplification
The moment hierarchy has direct consequences for sequential decision problems.
@@ 11.2 Strict value distortion
Proposition 11.1 (Exact value gap).
@@
Corollary 11.2 (Stopping boundary distortion).
@@
By Proposition 11.1, for each $\tau_{0}\in\{2,\ldots,K-1\}$, $\mathbb{E}[(1-\theta_{\infty})^{\tau_{0}}\mid\mathcal{F}_{n}]>(1-m_{n})^{\tau_{0}}$ when $\sigma_{n}^{2}>0$. Choose $r$ strictly between these two values; then $\tau_{0}$ lies in the Bayes feasible set but not the plug-in feasible set. ∎
The distortion $V_{n}-\widetilde{V}_{n}=\sigma_{n}^{2}$ vanishes at rate $O(1/n)$ as the posterior concentrates (Proposition 9.3), so predictive completeness is necessary for optimal sequential decisions.
@@ 12 Discussion
Three notions of inferential validity appear in the literature and bear on the results developed here. Blackwell admissibility requires that no other rule uniformly dominates a given rule; by Theorem 7.2, admissible rules are generalized Bayes. Anytime-valid inference controls type-I error at arbitrary stopping times via e-processes, nonnegative supermartingales under $H_{0}$ (Ville, 1939; Grünwald et al., 2024). Conformal validity achieves marginal coverage $\mathbb{P}(Y_{n+1}\in\hat{C}_{n})\geq 1-\alpha$ under exchangeability, finite-sample and distribution-free but set-valued. These notions address distinct inferential objectives, and multi-step predictives—polynomial functionals of $\Pi(\cdot\mid\mathcal{F}_{n})$—fall under the scope of Theorems 6.3 and 10.3.
For one-step prediction ($k=1$), first-moment coherence suffices: $\mathbb{P}(X_{n+1}=1\mid\mathcal{F}_{n})=\theta_{n}$.
For $k\geq 2$, the martingale condition alone does not uniquely identify $k$-step predictives (Theorem 6.3). A martingale posterior construction that specifies the full conditional law of $\theta_{\infty}\mid\mathcal{F}_{n}$ is predictively complete (Theorem 10.3); one that specifies only the update rule $\theta_{n}\mapsto\theta_{n+1}$ leaves multi-step predictives not uniquely determined, and the plug-in $(1-\theta_{n})^{k}$ is inadmissible for $k\geq 2$ (Corollary 7.4).
Many practical martingale-posterior constructions do in fact specify a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$, in which case the closure condition of Theorem 10.3 is automatically satisfied.
Example 12.1 (Diagnostic application).
@@
The moment hierarchy rests on the classical theory of moment problems (Hausdorff (1921); Feller (1971)). Hausdorff determinacy (that a measure on $[0,1]$ is uniquely determined by its moment sequence) is the key structural tool, used both to establish injectivity of the $k$-step run probabilities (Theorem 4.3) and to prove the closure theorem (Theorem 10.3). Compact support is essential: the moment problem is generically indeterminate for unbounded models (Remark 4.5), where analogous results remain open.
@@
Fong, Holmes, and Walker (2023) introduced the martingale posterior as a general-purpose alternative to likelihood-based updating. Their framework specifies a predictive resampling scheme that, in implementations, often amounts to specifying a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$. The insufficiency results of Section 6 apply to the abstract martingale condition, not to specific implementations that go beyond it. The closure theorem (Theorem 10.3) and the optimal stopping distortion (Section 11) together characterize the full consequences of first-moment versus full-law specification.
@@
Goldstein and Wooff (1998) provide a linear-Bayes treatment of adjusting exchangeable beliefs; their framework deliberately stops short of full distributional specification. Theorem 10.3 identifies the predictive consequences of that choice. The related contributions of Polson and Zantedeschi (2025) and Datta and Polson (2025) to finite-moment prediction and empirical Bayes are placed in context in the Introduction. Bayesian predictive inference via c.i.d. sequences, without a fully specified prior, is developed in Berti, Dreassi, Pratelli and Rigo (2021); Fortini and Petrone (2017) give a predictive characterization of exchangeable laws in a related setting. Leisen, Pratelli and Rigo (2025) establish that $X$ is c.i.d. if and only if $(\alpha_{n}(f))$ is a martingale for every bounded Borel $f$; the martingale posterior of Fong, Holmes and Walker (2023) constrains only $\alpha_{n}(\mathrm{id})=\theta_{n}$, and the present results quantify the predictive consequences of that gap through the moment hierarchy. In a complementary direction, Battiston and Cappello (2025) introduce almost conditional identically distributed (a.c.i.d.) sequences, in which predictive distributions are measure-valued almost supermartingales parametrized by a sequence of deviation parameters; where the present results identify what the strict martingale condition underdetermines, their framework broadens the class of admissible predictive schemes beyond the c.i.d. setting.
Several open questions emerge from this work.
@@
Minimal sufficient conditions. The closure theorem (Theorem 10.3) states that predictive completeness requires the full conditional law. A natural question is whether there exist finitely many moment constraints (beyond the first) that yield “approximate” coherence in a precise quantitative sense. The second-order bound (Proposition 6.4) implies that specifying $\mu_{1}$ and $\mu_{2}$ controls $k$-step predictives up to third-order remainders, but the precise minimax rates over the class of posteriors with given $(\mu_{1},\mu_{2})$ are unknown.
Beyond Bernoulli. The moment hierarchy exploits the compact support $[0,1]$ via the Hausdorff moment theorem. For exponential families with non-compact natural parameter space (Gaussian mean, Poisson rate), the moment problem is generically indeterminate (Remark 4.5). Whether analogous insufficiency results hold, and what replaces the Hausdorff determinacy, is open. Carleman’s condition provides a sufficient criterion in the one-dimensional case; the multivariate setting requires additional structure.
Computational implications. When predictive completeness is needed, one must specify the conditional law of $\theta_{\infty}\mid\mathcal{F}_{n}$. Whether approximate methods (variational, ABC) can be disciplined by moment constraints to achieve practical completeness is an open question.
Connection to calibration. Conformal prediction achieves marginal coverage without moment constraints (Table 4). Whether there exists a unified framework that interpolates between marginal calibration and full predictive coherence, parameterised by the number of constrained moments, would connect the conformal and Bayesian literatures.
@@
In summary, the martingale posterior framework provides a flexible mean-calibrated updating mechanism; the present results characterize the additional structure required for predictive completeness. Martingale coherence (1) constrains only the first conditional moment. Predictive completeness requires the full conditional law, equivalently all moments on $[0,1]$ (Theorem 10.3). The Bayes posterior supplies this; the martingale condition alone does not.
@@ Appendix B General functional non-identification
Theorem B.1 (General nonlinearity).
@@
The case $f(\theta)=(1-\theta)^{k}$ recovers Lemma 6.1. The case $f(\theta)=\mathbf{1}[\theta\leq t]$ shows that even the posterior CDF at a single point is not determined by the mean; the case $f(\theta)=-\theta\log\theta-(1-\theta)\log(1-\theta)$ (Bernoulli entropy, strictly concave) gives the same conclusion for expected entropy.

### 01 Apr 2026 v5

@@ 11 Optimal stopping and dynamic amplification
The moment hierarchy has direct consequences for sequential decision problems. Optimal stopping is the prototypical such problem: a clinician deciding when to stop a trial, a firm choosing when to exercise an option, or an analyst choosing when to “peek” at accumulating data must each evaluate multi-step continuation values. These values are polynomial functionals of the posterior and therefore depend on moments beyond the mean. The distortion introduced by ignoring higher moments is not merely theoretical: it shifts the optimal stopping boundary itself.
@@ 11.2 Strict value distortion
Proposition 48 (Exact value gap).
@@
Corollary 49 (Stopping boundary distortion).
@@
By Proposition 48, for each $\tau_{0}\in\{2,\ldots,K-1\}$, $\mathbb{E}[(1-\theta_{\infty})^{\tau_{0}}\mid\mathcal{F}_{n}]>(1-m_{n})^{\tau_{0}}$ when $\sigma_{n}^{2}>0$. Choose $r$ strictly between these two values; then $\tau_{0}$ lies in the Bayes feasible set but not the plug-in feasible set. ∎
The distortion $V_{n}-\widetilde{V}_{n}=\sigma_{n}^{2}$ vanishes at rate $O(1/n)$ as the posterior concentrates (Proposition 42), so predictive completeness is necessary for optimal sequential decisions.
@@ 12 Discussion
For one-step prediction ($k=1$), first-moment coherence suffices: $\mathbb{P}(X_{n+1}=1\mid\mathcal{F}_{n})=\theta_{n}$. For $k\geq 2$, the martingale condition alone does not uniquely identify $k$-step predictives (Theorem 23), and the plug-in $(1-\theta_{n})^{k}$ is inadmissible (Corollary 31). Predictive completeness requires the full conditional law of $\theta_{\infty}\mid\mathcal{F}_{n}$ (Theorem 47).
Many practical martingale-posterior constructions do in fact specify a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$, in which case the closure condition of Theorem 47 is automatically satisfied.
Example 50 (Diagnostic application).
@@
The moment hierarchy rests on the classical theory of moment problems (Hausdorff (1921); Feller (1971)). Hausdorff determinacy (that a measure on $[0,1]$ is uniquely determined by its moment sequence) is the key structural tool, used both to establish injectivity of the $k$-step run probabilities (Theorem 15) and to prove the closure theorem (Theorem 47). Compact support is essential: the moment problem is generically indeterminate for unbounded models (Remark 17), where analogous results remain open.
@@
Fong, Holmes, and Walker (2023) introduced the martingale posterior as a general-purpose alternative to likelihood-based updating. Their framework specifies a predictive resampling scheme that, in implementations, often amounts to specifying a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$. The insufficiency results of Section 6 apply to the abstract martingale condition, not to specific implementations that go beyond it. The closure theorem (Theorem 47) and the optimal stopping distortion (Section 11) together characterize the full consequences of first-moment versus full-law specification.
@@
Goldstein and Wooff (1998) provide a linear-Bayes treatment of adjusting exchangeable beliefs; their framework deliberately stops short of full distributional specification. Theorem 47 identifies the predictive consequences of that choice. The related contributions of Polson and Zantedeschi (2025) and Datta and Polson (2025) to finite-moment prediction and empirical Bayes are placed in context in the Introduction. Bayesian predictive inference via c.i.d. sequences, without a fully specified prior, is developed in Berti, Dreassi, Pratelli and Rigo (2021); Fortini and Petrone (2017) give a predictive characterization of exchangeable laws in a related setting. Leisen, Pratelli and Rigo (2026) characterize c.i.d. sequences via the martingale property of $(\alpha_{n}(f))$ for bounded Borel $f$. The martingale posterior of Fong, Holmes and Walker (2023) constrains only $\alpha_{n}(\mathrm{id})=\theta_{n}$; our results quantify the predictive consequences of that gap. In a complementary direction, Battiston and Cappello (2025) introduce almost conditional identically distributed (a.c.i.d.) sequences, in which predictive distributions are measure-valued almost supermartingales parametrized by a sequence of deviation parameters; where the present results identify what the strict martingale condition underdetermines, their framework broadens the class of admissible predictive schemes beyond the c.i.d. setting.
@@
Several questions remain open. First, whether finitely many moment constraints beyond the first yield “approximate” coherence in a precise quantitative sense: the second-order bound (Proposition 24) controls $k$-step predictives up to third-order remainders, but minimax rates over posteriors with given $(\mu_{1},\mu_{2})$ are unknown. Second, for exponential families with non-compact parameter space (Gaussian mean, Poisson rate), the moment problem is generically indeterminate (Remark 17), and what replaces Hausdorff determinacy is open. Third, whether approximate methods (variational, ABC) can be disciplined by moment constraints to achieve practical completeness.
@@
Martingale coherence (1) constrains only the first conditional moment. Predictive completeness requires the full conditional law, equivalently all moments on $[0,1]$ (Theorem 47). The Bayes posterior supplies this; the martingale condition alone does not.
Acknowledgements
We thank Luca Pratelli and Pietro Rigo for comments that helped reshape several sections of this paper.
@@ Appendix B General functional non-identification
Theorem 52 (General nonlinearity).
@@
The case $f(\theta)=(1-\theta)^{k}$ recovers Lemma 21. The case $f(\theta)=\mathbf{1}[\theta\leq t]$ shows that even the posterior CDF at a single point is not determined by the mean; the case $f(\theta)=-\theta\log\theta-(1-\theta)\log(1-\theta)$ (Bernoulli entropy, strictly concave) gives the same conclusion for expected entropy.